Circular No. 7812

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 C1 Word has been received of the independent visual discovery of a comet

by Kaoru Ikeya (Mori, Shuchi, Shizuoka, Japan; 0.25-m reflector, 39×; communicated by S. Nakano, Sumoto, Japan; coma diameter 2′ with weak condensation; motion ∼ 5′ northeastward in 30 min) and by Daqing Zhang (near Kaifeng, Henan province, China; 0.2-m reflector; communicated by J. Zhu, Peking University; coma diameter 3′).

2002 UT  α_{2000}  δ_{2000}  m_{1} Observer
Feb. 1.408 0^{h}08.^{m}9 −17^{o}42′ 9.0 Ikeya
1.47  0 09  −17 30  8.5 Zhang

V838 MONOCEROTIS = PECULIAR VARIABLE IN MONOCEROS

N. N. Samus, Institute of Astronomy, Moscow, informs us that the designation V838 Mon has been given to this unusual variable star (cf. IAUC 7786, 7789, 7791, 7796).

T. Zwitter, Department of Physics, University of Ljubljana; and U. Munari, Padova and Asiago Astronomical Observatories, report on CCD échelle spectroscopic observations (range 460–930 nm, resolution 18 000) of V838 Mon obtained on Jan. 26.0 UT with the Asiago 1.82-m telescope. The spectrum is that of a heavily reddened cool giant K-type star at heliocentric velocity +53 km/s. Deep P-Cyg profiles mark the Ca II triplet at 849–866 nm, S I (869.3 and 675.7 nm), Si I (614.2 nm), Na I (589.0 and 589.6 nm), Ca I (650.0, 585.7, and 527.0 nm), Ni I (677.2 nm), Y II (566.3 nm), Zr II (611.5 nm), and C I (661.1 and 667.2 nm), as well as many Cr I, Gd II, Fe I, and Ti I (but not Ba II) lines. The average terminal heliocentric absorption velocity in the P-Cyg profile is −458 km/s, and the minimum is at −252 km/s. Hα appears as a narrow absorption at +54 km/s (FWHM 41 km/s, equivalent width 0.064 nm), superimposed on a marginal emission at +72 km/s. Interstellar Na I and K I lines have two components at +47 and +73 km/s with component intensities corresponding to E(B − V) = 0.56 and 0.24, respectively, using the calibration of Munari and Zwitter (1997, A.Ap. 318, 269). The total extinction E(B − V) = 0.80 is confirmed by the 0.029-nm equivalent width of the diffuse interstellar band at 862.1 nm. The reddening suggests a distance > 3 kpc, judging from extinction maps of Neckel and Klare (1980, A.Ap. Suppl. 42, 251).

2002 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green